# Splitting and gluing in sine-dilaton gravity: matter correlators and the wormhole Hilbert space

### v1

@@ 3.2 Multi-particle Hilbert space
We note again that in our multi-particle Hilbert space, the energy eigenstates exhibit a non-trivial state-dependent structure, while the length eigenstates simply factorize. This is a direct feature that follows from our splitting and gluing procedure in sine-dilaton gravity. The quasi-unitary transformation operator between these two bases has the same form as in (83), which is non-local between different regions separated by the matter lines. This is in contrast to previous constructions by Lin and Stanford [15, 16], which focus on the non-factorized length (or chord number) basis. [^15] For a more careful comparison, check footnote 14. It would be interesting to find an isomorphism between our construction with theirs, possibly similar to the one discussed in [28].
@@ 4.1 OTOC in 0-split representation
$=\sum_{n=0}^{\infty}\frac{q^{2n\Delta^{\prime}}}{(q^{2},q^{4\Delta};q^{2})_{n}}\leavevmode\nobreak\ Q_{n}(\cos(\theta_{3})|q^{2\Delta}e^{i\theta_{4}},q^{2\Delta}e^{-i\theta_{4}};q^{2})\leavevmode\nobreak\ Q_{n}(\cos(\theta_{2})|q^{2\Delta}e^{i\theta_{1}},q^{2\Delta}e^{-i\theta_{1}};q^{2})$
@@
$\prod_{i=1}^{4}\left(\int_{0}^{\pi}d\theta_{i}\exp\left(\beta_{i}\frac{\cos(\theta_{i})}{\hbar}\right)\rho(\theta_{i})\right)\leavevmode\nobreak\ R_{\theta_{2}\theta_{4}}^{q^{2}}\begin{bmatrix}\theta_{1}&\Delta\\ \theta_{3}&\Delta^{\prime}\end{bmatrix}$
@@ 4.2 OTOC in 1-split representation
$=R_{\theta_{2}\theta_{4}}^{q^{2}}\begin{bmatrix}\theta_{1}&\Delta_{L}+\Delta_{R}\\ \theta_{3}&\Delta^{\prime}\end{bmatrix}\leavevmode\nobreak\ \frac{(q^{4(\Delta_{L}+\Delta_{R})};q^{2})_{\infty}}{[(q^{2(\Delta_{L}+\Delta_{R}))}e^{i(\pm\theta_{1}\pm\theta_{2})},q^{2(\Delta_{L}+\Delta_{R}))}e^{i(\pm\theta_{3}\pm\theta_{4})};q^{2})_{\infty}]^{1/2}}.$ (116)
@@
$\prod_{i=1}^{4}\left(\int_{0}^{\pi}d\theta_{i}\exp\left(\beta_{i}\frac{\cos(\theta_{i})}{\hbar}\right)\rho(\theta_{i})\right)\leavevmode\nobreak\ R_{\theta_{2}\theta_{4}}^{q^{2}}\begin{bmatrix}\theta_{1}&\Delta\\ \theta_{3}&\Delta^{\prime}\end{bmatrix}$
@@
The formula (116) is a new identity for the 6j-symbols161616 To our knowledge, this identity does not appear in the existing literature. It may, however, be related to the Biedenharn-Eliott identity of quantum group $\mathcal{U}_{q}(su(1,1))$ [40]. We thank Misha Isachenkov for pointing this out., which is similar to the familiar ”pentagon identity”, but with the ends of the vertical lines attached together. We present this relation pictorially
$\int_{0}^{\pi}\int_{0}^{\pi}d\bar{\alpha}_{1}d\bar{\alpha}_{2}\,\rho(\bar{\alpha}_{1})\,\rho(\bar{\alpha}_{2})\quad\leavevmode\hbox{$\vbox{\hbox{\resizebox{}{}{{\leavevmode\hbox{\set@color{}}}}}}$}=\quad\leavevmode\hbox{$\vbox{\hbox{\resizebox{}{}{{\leavevmode\hbox{\set@color{}}}}}}$}.$ (118)
@@ 5 General n-point functions and the Feynman rules
$=$ $\prod_{i=1}^{6}\left(\int_{0}^{\pi}d\theta_{i}\exp\left(\beta_{i}\frac{\cos(\theta_{i})}{\hbar}\right)\rho(\theta_{i})\right)\leavevmode\nobreak\ R_{\theta_{2}\theta_{6}}^{q^{2}}\begin{bmatrix}\theta_{1}&\Delta_{1}\\ \theta_{5}&\Delta_{3}\end{bmatrix}\,R_{\theta_{3}\theta_{5}}^{q^{2}}\begin{bmatrix}\theta_{2}&\Delta_{2}\\ \theta_{4}&\Delta_{3}\end{bmatrix}$
@@
Note that the bulk two-point function $\langle\theta_{5}|e^{-\Delta_{3}\hat{L}_{2}}|\theta_{2}\rangle$ is canceled by the factors from other two-point functions in (119). The result involves the convolution of two R-matrices. However, one cannot use (116) to reduce it to a single R-matrix due to the extra $\exp\left({\beta_{i}\frac{\cos(\theta_{i})}{\hbar}}\right)$ factors.
@@ 5.3 Feynman rules
    $\leavevmode\hbox{$\vbox{\hbox{\resizebox{}{}{{\leavevmode\hbox{\set@color{}}}}}}$}=\left(\frac{(q^{4\Delta};q^{2})_{\infty}}{(q^{2\Delta}e^{i(\pm\theta_{i}\pm\theta_{j})};q^{2})_{\infty}}\right)^{1/2}.$ (131)
@@
    $\leavevmode\hbox{$\vbox{\hbox{\resizebox{}{}{{\leavevmode\hbox{\set@color{}}}}}}$}=\exp\left(\beta_{i}\frac{\cos(\theta_{i})}{\hbar}\right).$ (132)
@@
    $\leavevmode\hbox{$\vbox{\hbox{\resizebox{}{}{{\leavevmode\hbox{\set@color{}}}}}}$}=R_{\theta_{j}\theta_{\ell}}^{q^{2}}\begin{bmatrix}\theta_{i}&\Delta_{2}\\ \theta_{k}&\Delta_{1}\end{bmatrix}$. (133)
@@
    $\leavevmode\hbox{$\vbox{\hbox{\resizebox{}{}{{\leavevmode\hbox{\set@color{}}}}}}$}=1.$ (134)
@@ 6 JT gravity limit
Our splitting and gluing process is not directly avialable in JT gravity, as it uses essentially the description of EOW branes by two parameters, only one of which is visible in the usual EOW brane formulations of JT gravity [37]. Nevertheless, the resulting formulas do have a good JT limit, which we now describe briefly.
@@
In this limit, most quantities in sine-dilaton gravity diverges, and they map to their JT gravity counterparts as follows
@@ 7.1 Concrete open questions
  • Our multi-particle Hilbert space enables the directly bulk computation of most correlation functions, except for those involving the Youg-Baxter move symmetry. For example:
@@
    This diagram is expected to follow the proposed Feynman rules, but a direct bulk calculation would require handling the crossing of matter lines in $n$-particle states. While such crossings should of course just reproduce a 6j-symbol, making this explicit within our splitting and gluing procedure might require extra work.
@@ 7.2 Future directions
Another fascinating future directions is a better understanding of purely bulk regions that are disconnected from the boundary, using our multi-particle Hilbert space. In some way the state-dependence structure we find is to be expected from previous studies of bulk reconstruction and quantum error correction. However, we expect a new layer of non-locality or non-commutativity to emerge due to the quantum group structure. By comparing to the JT limit, it may be possible to disentangle these two aspects of bulk reconstruction.
@@ Acknowledgments
The authors thank Micha Berkooz, Andreas Blommaert, Misha Isachenkov, Thomas Mertens and Abhisek Sahu for interesting discussions. The work is supported by a Discovery grant from NSERC.
@@ Appendix A Pizza slicing representation of the OTOC
$\prod_{i=1}^{5}\left(\int_{0}^{\pi}d\bar{\alpha}_{i}\,\rho(\bar{\alpha}_{i})\right)\leavevmode\hbox{$\vbox{\hbox{\resizebox{}{}{{\leavevmode\hbox{\set@color{}}}}}}$}=\,\,\,\leavevmode\hbox{$\vbox{\hbox{\resizebox{}{}{{\leavevmode\hbox{\set@color{}}}}}}$}$, (151)

### v2

@@ 3.2 Multi-particle Hilbert space
We note again that in our multi-particle Hilbert space, the energy eigenstates exhibit a non-trivial state-dependent structure, while the length eigenstates simply factorize. This is a direct feature that follows from our splitting and gluing procedure in sine-dilaton gravity. The quasi-unitary transformation operator between these two bases has the same form as in (83), which is non-local between different regions separated by the matter lines. This is in contrast to previous constructions by Lin and Stanford [15, 16], which focus on the non-factorized length (or chord number) basis. [^17] For a more careful comparison, check footnote 16. It would be interesting to find an isomorphism between our construction with theirs, possibly similar to the one discussed in [30].
@@ 4.1 OTOC in 0-split representation
$=\sum_{n=0}^{\infty}\frac{q^{2n\Delta^{\prime}}}{(q^{2},q^{4\Delta};q^{2})_{n}}\penalty 10000\ Q_{n}(\cos(\theta_{3})|q^{2\Delta}e^{i\theta_{4}},q^{2\Delta}e^{-i\theta_{4}};q^{2})\penalty 10000\ Q_{n}(\cos(\theta_{2})|q^{2\Delta}e^{i\theta_{1}},q^{2\Delta}e^{-i\theta_{1}};q^{2})$
@@
$\prod_{i=1}^{4}\left(\int_{0}^{\pi}d\theta_{i}\exp\left(\beta_{i}\frac{\cos(\theta_{i})}{\hbar}\right)\rho(\theta_{i})\right)\penalty 10000\ R_{\theta_{2}\theta_{4}}^{q^{2}}\begin{bmatrix}\theta_{1}&\Delta\\ \theta_{3}&\Delta^{\prime}\end{bmatrix}$
@@ 4.2 OTOC in 1-split representation
$=R_{\theta_{2}\theta_{4}}^{q^{2}}\begin{bmatrix}\theta_{1}&\Delta_{L}+\Delta_{R}\\ \theta_{3}&\Delta^{\prime}\end{bmatrix}\penalty 10000\ \frac{(q^{4(\Delta_{L}+\Delta_{R})};q^{2})_{\infty}}{[(q^{2(\Delta_{L}+\Delta_{R}))}e^{i(\pm\theta_{1}\pm\theta_{2})},q^{2(\Delta_{L}+\Delta_{R}))}e^{i(\pm\theta_{3}\pm\theta_{4})};q^{2})_{\infty}]^{1/2}}.$ (116)
@@
$\prod_{i=1}^{4}\left(\int_{0}^{\pi}d\theta_{i}\exp\left(\beta_{i}\frac{\cos(\theta_{i})}{\hbar}\right)\rho(\theta_{i})\right)\penalty 10000\ R_{\theta_{2}\theta_{4}}^{q^{2}}\begin{bmatrix}\theta_{1}&\Delta\\ \theta_{3}&\Delta^{\prime}\end{bmatrix}$
@@
The formula (116) is a new identity for the 6j-symbols [^18] To our knowledge, this identity does not appear in the existing literature. It may, however, be related to the Biedenharn-Eliott identity of quantum group $\mathcal{U}_{q}(su(1,1))$ [44]. We thank Misha Isachenkov for pointing this out., which is similar to the familiar ”pentagon identity”, but with the ends of the vertical lines attached together. We present this relation pictorially
$\int_{0}^{\pi}\int_{0}^{\pi}d\bar{\alpha}_{1}d\bar{\alpha}_{2}\,\rho(\bar{\alpha}_{1})\,\rho(\bar{\alpha}_{2})\quad\hbox{$\vbox{\hbox{\resizebox{}{}{{\hbox{\set@color{}}}}}}$}=\quad\hbox{$\vbox{\hbox{\resizebox{}{}{{\hbox{\set@color{}}}}}}$}.$ (118)
@@ 5 General n-point functions and the Feynman rules
$=$ $\prod_{i=1}^{6}\left(\int_{0}^{\pi}d\theta_{i}\exp\left(\beta_{i}\frac{\cos(\theta_{i})}{\hbar}\right)\rho(\theta_{i})\right)\penalty 10000\ R_{\theta_{2}\theta_{6}}^{q^{2}}\begin{bmatrix}\theta_{1}&\Delta_{1}\\ \theta_{5}&\Delta_{3}\end{bmatrix}\,R_{\theta_{3}\theta_{5}}^{q^{2}}\begin{bmatrix}\theta_{2}&\Delta_{2}\\ \theta_{4}&\Delta_{3}\end{bmatrix}$
@@
Note that the bulk two-point function $\langle\theta_{5}|e^{-\Delta_{3}\hat{L}_{2}}|\theta_{2}\rangle$ is canceled by the factors from other two-point functions in (119), in agreement with [45]. The result involves the convolution of two R-matrices. However, one cannot use (116) to reduce it to a single R-matrix due to the extra $\exp\left({\beta_{i}\frac{\cos(\theta_{i})}{\hbar}}\right)$ factors.
@@ 5.3 Feynman rules
    $\hbox{$\vbox{\hbox{\resizebox{}{}{{\hbox{\set@color{}}}}}}$}=\left(\frac{(q^{4\Delta};q^{2})_{\infty}}{(q^{2\Delta}e^{i(\pm\theta_{i}\pm\theta_{j})};q^{2})_{\infty}}\right)^{1/2}.$ (131)
@@
    $\hbox{$\vbox{\hbox{\resizebox{}{}{{\hbox{\set@color{}}}}}}$}=\exp\left(\beta_{i}\frac{\cos(\theta_{i})}{\hbar}\right).$ (132)
@@
    $\hbox{$\vbox{\hbox{\resizebox{}{}{{\hbox{\set@color{}}}}}}$}=R_{\theta_{j}\theta_{\ell}}^{q^{2}}\begin{bmatrix}\theta_{i}&\Delta_{2}\\ \theta_{k}&\Delta_{1}\end{bmatrix}$. (133)
@@
    $\hbox{$\vbox{\hbox{\resizebox{}{}{{\hbox{\set@color{}}}}}}$}=1.$ (134)
@@ 6 JT gravity limit
Our splitting and gluing process is not directly available in JT gravity, as it uses essentially the description of EOW branes by two parameters, only one of which is visible in the usual EOW brane formulations of JT gravity [40]. Nevertheless, the resulting formulas do have a good JT limit, which we now describe briefly.
@@
In this limit, most quantities in sine-dilaton gravity diverge, and they map to their JT gravity counterparts as follows
@@ 7.1 Concrete open questions
  • Our multi-particle Hilbert space enables the directly bulk computation of most correlation functions, except for those involving the Yang-Baxter move symmetry. For example:
@@
    This diagram is expected to follow the proposed Feynman rules, but a direct bulk calculation would require handling the crossing of matter lines in $n$-particle states. While such crossings should of course just reproduce a 6j-symbol, making this explicit within our splitting and gluing procedure might require extra work. [^25] A concrete obstruction is that, in this case, three vertices form a triangular region in the center of the bulk. Once one line in the bulk is treated as a vertical EOW brane, the calculation of two vertices on this line automatically fixes the remaining two lines to be horizontal operators. As a result, there is inevitably at least one vertex at which two lines are both horizontal. For such a vertex, our splitting–gluing procedure does not provide a well-defined prescription, and therefore cannot be directly applied to compute the corresponding contribution.
@@ 7.2 Future directions
Note added: The quantum group used in [33] is $\mathcal{U}_{q}(sl(2,\mathbb{R}))$. However, after our paper appeared, a later work [52] reproduced exactly the same multi-particle Hilbert space structure as ours using the coproduct of the quantum group $\mathcal{U}_{q}(su(1,1))$. This suggests that the quantum group $\mathcal{U}_{q}(su(1,1))$ is more directly relevant to our construction.
@@
Another fascinating future direction is a better understanding of purely bulk regions that are disconnected from the boundary, using our multi-particle Hilbert space. In some way the state-dependence structure we find is to be expected from previous studies of bulk reconstruction and quantum error correction. However, we expect a new layer of non-locality or non-commutativity to emerge due to the quantum group structure. By comparing to the JT limit, it may be possible to disentangle these two aspects of bulk reconstruction.
@@ Acknowledgments
The authors thank Sergio Aguilar, Micha Berkooz, Andreas Blommaert, Hong Zhe (Vincent) Chen, Misha Isachenkov, Thomas Mertens, Abhisek Sahu and Jiuci Xu for interesting discussions. The work is supported by a Discovery grant from NSERC.
@@ Appendix A Pizza slicing representation of the OTOC
$\prod_{i=1}^{5}\left(\int_{0}^{\pi}d\bar{\alpha}_{i}\,\rho(\bar{\alpha}_{i})\right)\hbox{$\vbox{\hbox{\resizebox{}{}{{\hbox{\set@color{}}}}}}$}=\,\,\,\hbox{$\vbox{\hbox{\resizebox{}{}{{\hbox{\set@color{}}}}}}$}$, (151)